# Benchmarking Universal Machine Learning Interatomic Potentials for Elastic Property Prediction

### v2

@@ II.3 Elastic Property Calculations
For structure optimization and elastic-property calculations, we employ the Atomic Simulation Environment (ASE) [ase-paper, ISI:000175131400009] and Pymatgen [ong2013python]. The FIRE algorithm [bitzek2006structural] is used for energy minimization and the FretchCellFilter [ase-paper] is applied to preserve space group symmetry during relaxation. The force convergence criterion was set to 0.1 eV/Åfor structure relaxation, ensuring mechanical equilibrium before strain application.
Once the elastic tensor is obtained, the bulk modulus, shear modulus, Young’s modulus, Poisson’s ratio and other derived mechanical properties can be calculated. In this work, all derived mechanical properties are Voigt-Reuss-Hill average values via MechElastic [singh2021mechelastic] analysis module. The Young’s modulus $E$ and Poisson’s ratio $\nu$ are obtained based on the bulk modulus $K$ and shear modulus $G$ as follows:
@@
III Results and Analysis
III.1 Model Performance Analysis
@@
Figure 3: Distributions of (a) bulk modulus, (b) shear modulus, (c) Young’s modulus, and (d) Poisson’s ratio, computed as Voigt–Reuss–Hill (VRH) averages from DFT and the four uMLIP models. Each violin plot shows the overall distribution, with blue and red dashed lines marking the median and mean, and short lines denoting the extrema.
As shown in Fig. 3, we compare the distributions of bulk, shear, and Young’s moduli and Poisson’s ratio obtained from DFT and from the four uMLIP models. Throughout this paragraph, values given in parentheses denote mean model predictions. Overall, all models reproduce the macroscopic DFT trends, but systematic deviations remain in their absolute magnitudes. For the bulk modulus, the DFT mean is 104.1 GPa, while the model means lie in the range 100–115 GPa, indicating robust performance in capturing volumetric compressibility. In contrast, discrepancies are more pronounced for the shear and Young’s moduli: the DFT means are 47.3 and 122.5 GPa, whereas CHGNet yields the lowest values (28.6 and 77.5 GPa), systematically underestimating rigidity; MACE (58.2 and 148.1 GPa) and SevenNet (56.3 and 143.9 GPa) overestimate both moduli, reflecting a tendency to over-enhance stiffness; MatterSim (50.8 and 130.5 GPa) gives intermediate results that remain closest to the DFT benchmarks. For Poisson’s ratio, the DFT mean is 0.291: CHGNet substantially overestimates it (0.371), implying artificially high ductility, MACE and SevenNet slightly underestimate it (0.279 and 0.282), and MatterSim (0.294) is nearly indistinguishable from DFT. Taken together, these distributions indicate that, while all models capture the overall elastic trends, notable systematic biases persist, particularly for the shear and Young’s moduli and for Poisson’s ratio.
@@
To enable quantitative evaluation, we present point-wise comparisons of the primary elastic properties in Fig. 4. For the bulk modulus, SevenNet and MACE exhibit the highest consistency with DFT, achieving correlation coefficients of approximately $R\approx 0.94$ and mean absolute errors (MAE) around 15 GPa, outperforming both CHGNet ($R=0.909$) and MatterSim ($R=0.924$). For the shear modulus, MACE attains the highest correlation ($R=0.896$), followed by SevenNet ($R=0.895$), while MatterSim yields intermediate accuracy ($R=0.847$) and CHGNet remains significantly weaker ($R=0.546$). Regarding the Young’s modulus, MatterSim yields the mean closest to DFT, but in this correlation-centric assessment MACE attains the higher correlation (R = 0.901) than MatterSim (R = 0.860); SevenNet is lower (R = 0.791), and CHGNet remains the weakest (R = 0.546). For the Poisson’s ratio, a different trend emerges: MACE and MatterSim achieve significantly higher correlations ($R\approx 0.65$) than CHGNet ($R=0.301$) and SevenNet ($R=0.374$), indicating their robustness in capturing ratio-type properties. Overall, MACE and SevenNet alternate in leading performance depending on the property considered, while MatterSim also exhibits consistently reliable behavior, achieving mean values closest to DFT and competitive correlations across most properties. The relative superiority of these models remains task-dependent, reflecting the varying accuracy of current universal machine-learning interatomic potentials across different elastic property regimes.
@@
Beyond elasticity, stability classification provides another essential benchmark for evaluating model performance. Fig. 5 compares the stability predictions of the four uMLIPs against DFT references. SevenNet and MACE achieve the highest performance, with accuracies of 98.3% and 98.1%, respectively, and F1 scores approaching 0.99, reflecting well-balanced capability in identifying both stable and unstable materials. MatterSim ranks closely behind, while CHGNet reaches only 93.4% accuracy, significantly lower than the others and showing a higher rate of missed unstable samples. The confusion matrix analysis further indicates that both MACE and SevenNet exhibit consistently high precision ($\approx 0.997$) and recall (¿0.98), underscoring their robustness and reliability in large-scale stability screening tasks. Cross-model agreement analysis indicates strong overlap, with more than 10,700 materials consistently classified by both MACE and SevenNet in line with DFT. This highlights their superior generalization ability in stability classification across diverse material systems.
@@
III.2 Systematic Error Analysis
@@
Fig. 6 presents the relative error distributions of the four uMLIPs with respect to DFT values across various elastic descriptors and the values in parentheses in this paragraph correspond to median relative errors. CHGNet exhibits pronounced systematic deviations across most properties. In bulk modulus predictions, it shows a median error of $2.61\%$, indicating a tendency toward underestimation, whereas MACE and SevenNet display slight overestimations ($4.16\%$ and $2.88\%$, respectively), and MatterSim remains close to zero bias ($1.98\%$). For the shear modulus and Young’s modulus, CHGNet strongly underestimates both ($48.02\%$ and $44.20\%$), in sharp contrast to the overestimations observed for MACE ($13.83\%$ and $12.43\%$) and SevenNet ($9.79\%$ and $8.89\%$), while MatterSim again yields nearly symmetric distributions ($2.12\%$ and $2.24\%$). For Poisson’s ratio, CHGNet systematically overestimates ($27.25\%$), opposite to the mild underestimations of MACE ($4.35\%$) and SevenNet ($3.40\%$), whereas MatterSim remains almost unbiased ($0.70\%$). The bulk/shear ratio further highlights CHGNet’s strong positive bias ($77.05\%$), while the other models show values close to zero. CHGNet also shows especially high variability in anisotropy metrics, suggesting instability in capturing complex anisotropic behavior. For Cauchy pressure, CHGNet exhibits a systematic positive bias, whereas the others lean toward negative deviations. The large deviations in the predicted anisotropy and Cauchy pressure mainly reflect the high sensitivity of these quantities to small differences among elastic constants. In this work, most materials exhibit a relatively small degree of elastic anisotropy, with a DFT average of 1.97. The Cauchy pressure, being a difference quantity defined as $(C_{12}-C_{44})$, has a DFT average value of 17.9 GPa, which is much smaller than the bulk and Young’s modulus that are generally on the order of 100 GPa. Consequently, even moderate relative errors in the stiffness components can lead to large percentage deviations in these derived quantities, indicating that further development of uMLIPs to improve their accuracy in elastic property predictions is essential. Finally, in Debye temperature predictions, CHGNet again underestimates ($25.89\%$), while MACE ($6.55\%$) and SevenNet ($4.89\%$) perform closer to DFT, and MatterSim achieves the most balanced performance ($0.69\%$). Overall, CHGNet displays consistent systematic biases across multiple properties, whereas MACE, MatterSim, and SevenNet yield more symmetric, near-zero error distributions, reflecting higher robustness.
@@
To provide a clearer comparison of overall performance, Fig. 7 summarizes the MAPE values of different models across all elastic properties. It is evident that CHGNet systematically yields the highest error levels, with an average MAPE of 71.8%, underscoring its structural deficiencies in elastic property prediction. In contrast, SevenNet consistently achieves the lowest error, with an average MAPE of only 27.53%, highlighting its superior overall accuracy. Further analysis reveals that differences among models are relatively small for bulk and Young’s modulus, whereas much larger discrepancies arise in the bulk/shear ratio, universal anisotropy index and Cauchy pressure—properties closely linked to mechanical stability and anisotropy. Particularly noteworthy is that CHGNet’s MAPE exceeds 90% for these metrics, reflecting structural limitations in capturing the complex couplings within elastic tensors. While MACE and MatterSim outperform CHGNet, their overall accuracy remains inferior to SevenNet, reinforcing the conclusion that SevenNet exhibits stronger generalization capability in modeling the nonlinear interdependencies among elastic properties.
@@
III.3 Fundamental Limitations and the Impact of Fine-Tuning
@@
The limited performance of current uMLIPs in elastic property prediction can be partially attributed to training datasets dominated by equilibrium configurations, which do not provide sufficient coverage of strained states that are important for learning accurate mechanical responses. Fine-tuning [Fine-Tuning] provides a practical route to mitigating such systematic biases. To evaluate this strategy, we constructed a targeted fine-tuning dataset comprising 185 materials with the largest baseline errors, together with DFT-computed energies of their deformed structures, thereby explicitly introducing non-equilibrium information into the training domain. The DFT calculation settings are detailed in the Supporting Information. Fine-tuning on this targeted dataset yields well-controlled training mean absolute errors (MAEs): CHGNet achieved energy/force/stress MAEs of 24.089 meV/atom, 26.831 meV/Å, and 1.960 meV/Å3; MACE obtained 6.17 meV/atom, 26.38 meV/Å, and 5.67 meV/Å3; MatterSim reached 17.89 meV/atom, 9.574 meV/Å, and 0.836 meV/Å3; and SevenNet yielded 1.692 meV/atom, 5.101 meV/Å, and 0.528 meV/Å3.
All four uMLIPs—CHGNet, MACE, MatterSim, and SevenNet—were fine-tuned on this dataset. The updated potentials were then reassessed on the same 185 materials. Fig. S2 and Fig. S3 display heatmaps of the MAPE before and after fine-tuning, while Fig. S4 and Fig. S5 present the corresponding distributions of relative errors. To quantify these effects, Table 1 summarizes the relative change in MAPE, defined as the difference between the fine-tuned MAPE and the original MAPE, normalized by the original value. CHGNet shows the largest improvement, with an average MAPE reduction of 23.2%, followed by MatterSim at 20.7% and SevenNet at 18.0%. In contrast, MACE’s average MAPE increases by 13.8%, indicating a less favorable response to the added deformation information. Pronounced improvements occur for the two dimensionless ratios—the Poisson’s ratio and the bulk-to-shear modulus ratio—whose MAPE reductions exceed 40% for CHGNet, MatterSim, and SevenNet. In comparison, the six dimensional elastic constants (e.g., bulk modulus, shear modulus, Young’s modulus) typically show more modest reductions around 10–20%, highlighting a differential sensitivity to the inclusion of strained configurations. Table 2 reports the corresponding interquartile ranges (IQR), showing that fine-tuning leads to IQR reductions in seven of the eight elastic properties for CHGNet, and in all eight properties for both MatterSim and SevenNet. By contrast, MACE shows IQR reduction in only two properties and increased dispersion in the remaining six. These results demonstrate a clear difference in fine-tuning consistency among the four uMLIPs. For CHGNet, MatterSim, and SevenNet, the simultaneous improvements in both MAPE and IQR across all eight mechanical quantities indicate a coherent propagation of the energy–force–stress refinements into uniformly improved elastic responses. This reflects a high degree of physical consistency, where the optimization of the fundamental quantities translates directly into better predictions of derived mechanical observables. In contrast, the inconsistent behavior of MACE—improving in only a subset of properties while degrading in others—suggests weaker coupling between its learned potential-energy surface and its strain-dependent responses, pointing to limited fine-tuning consistency and reduced robustness to strained configurations. Such behavior illustrates that fine-tuning not only reduces systematic bias for CHGNet, MatterSim, and SevenNet but also yields more stable and concentrated error distributions. The degradation in MACE’s performance may reflect overfitting or architectural sensitivities that warrant further investigation.
These findings confirm that targeted fine-tuning with non-equilibrium configurations provides an effective and physically consistent route for improving the mechanical predictive capability of CHGNet, MatterSim, and SevenNet, whereas MACE shows limited and inconsistent gains.
@@
IV Discussion
IV.1 Implications for Materials Design Applications
@@
IV.2 Future Directions
As demonstrated in our fine-tuning study, incorporating strained configurations into the training process can effectively improve the predictive accuracy of uMLIPs for elastic properties, underscoring the importance of dataset diversity. Future developments should focus on systematically incorporating deformed structures into training datasets, for instance through active learning strategies that aim to improve mechanical property accuracy [zhang2020dp]. For elastic property predictions within specific chemical spaces, constructing domain-specific fine-tuning [deng2025systematic] datasets and adapting pretrained models accordingly could effectively mitigate systematic biases in those regions.
@@ IV.2 Future Directions
V Conclusions
Our benchmark study establishes the first systematic evaluation framework for applying uMLIPs to elastic property prediction, validated across nearly 11,000 crystalline materials. The results demonstrate clear differences in model suitability: SevenNet delivers the highest overall accuracy, MatterSim and MACE achieve a favorable balance between accuracy and computational efficiency, while CHGNet, constrained by its architectural design, performs relatively less effectively. These results provide quantitative, evidence-based guidance for selecting suitable uMLIPs in mechanical property calculations, ensuring that model performance is properly matched to application requirements.
@@ Acknowledgements.
Declarations
During manuscript preparation, we utilized GPT-5 LLM to enhance sentence structure, readability, and coherence. We also used Claude 4 LLM for optimizing the Python plotting and simulation scripts.
@@
All 10,994 crystal structures and the corresponding DFT-computed elastic property data used in this study are publicly available through the Materials Project database (https://materialsproject.org).
@@
Appendix A Supporting Information
The Supporting Information provides DFT computational details together with supplementary benchmarking analyses for the four uMLIPs evaluated in this work.
DFT Computational Details
All deformed configurations were computed using density functional theory (DFT) [kohn1965self] within the Vienna Ab initio Simulation Package (VASP) [VASP, VASP2]. The Perdew–Burke–Ernzerhof (PBE) generalized gradient approximation [PBE] and projector augmented-wave (PAW) [PAW] pseudopotentials were employed. To ensure strict consistency with the standardized Materials Project workflow, we adopted the same computational parameters: a plane-wave cutoff energy of 520 eV, electronic convergence set to $10^{-5}$ eV, and a $\Gamma$-centered Monkhorst–Pack $k$-point mesh [MP] generated with 1000/(number of atoms per cell) $k$-point. Spin polarization was enabled for all materials. For each strained configuration, a single-point DFT calculation was carried out to obtain total energies, forces, and stresses for use in fine-tuning the uMLIPs.
Supplementary Figures
Fig. S1 compares the processing-time distributions of the four uMLIPs for elastic-property calculations. Figs. S2 and S3 compare the mean absolute percentage errors for eight elastic properties before and after fine-tuning. Figs. S4 and S5 present the corresponding relative-error distributions.
Figure S1: Processing time distribution comparison across uMLIPs for elastic-property calculations. Each boxplot shows the median (solid line), mean (dashed line), and interquartile range (colored box), with short lines marking the extrema of the non-outlier data. Outliers were filtered using a z-score method (factor = 3.0) during preprocessing and are not displayed.
Figure S2: MAPE heatmap for elastic properties predicted by CHGNet, MACE, MatterSim, and SevenNet relative to DFT reference values before fine-tuning, for the selected 185 materials with high prediction errors. The properties analyzed include bulk modulus, shear modulus, Young’s modulus, Poisson’s ratio, bulk/shear ratio, universal anisotropy index, Cauchy pressure, and Debye temperature. Darker colors indicate higher errors, with values annotated in each cell.
Figure S3: MAPE heatmap for elastic properties predicted by CHGNet, MACE, MatterSim, and SevenNet relative to DFT reference values after fine-tuning, for the selected 185 materials with high prediction errors. The properties analyzed include bulk modulus, shear modulus, Young’s modulus, Poisson’s ratio, bulk/shear ratio, universal anisotropy index, Cauchy pressure, and Debye temperature. Darker colors indicate higher errors, with values annotated in each cell.
Figure S4: (a)-(h) Distribution of relative errors (%) for CHGNet, MACE, MatterSim, and SevenNet compared with DFT reference values across eight elastic properties for the selected 185 materials with high prediction errors before fine-tuning. Each boxplot shows the median (blue dashed line), mean (red dashed line), interquartile range (colored box), and the overall error range (short lines marking the extrema), with outliers omitted for clarity. The horizontal dashed line denotes zero error.
Figure S5: (a)-(h) Distribution of relative errors (%) for CHGNet, MACE, MatterSim, and SevenNet compared with DFT reference values across eight elastic properties for the selected 185 materials with high prediction errors after fine-tuning. Each boxplot shows the median (blue dashed line), mean (red dashed line), interquartile range (colored box), and the overall error range (short lines marking the extrema), with outliers omitted for clarity. The horizontal dashed line denotes zero error.

### v3

@@ 2.3 Elastic Property Calculations
For structure optimization and elastic-property calculations, we employ the Atomic Simulation Environment (ASE) [40, 41] and Pymatgen [42]. The fast inertial relaxation engine (FIRE) algorithm [43] is used for energy minimization and the FretchCellFilter [40] is applied to preserve space group symmetry during relaxation. The force convergence criterion was set to 0.1 eV/Å for structure relaxation, ensuring mechanical equilibrium before strain application.
The elastic tensor was calculated using a finite-strain approach as implemented in the MatCalc workflow [44]. Once the elastic tensor is obtained, the bulk modulus, shear modulus, Young’s modulus, Poisson’s ratio and other derived mechanical properties can be calculated. In this work, all derived mechanical properties are reported as Voigt–Reuss–Hill (VRH) averaged values, obtained using the MechElastic [45] analysis module. The Young’s modulus $E$ and Poisson’s ratio $\nu$ are obtained based on the bulk modulus $K$ and shear modulus $G$ as follows:
@@
3 Results and Analysis
3.1 Model Performance Analysis
@@
Figure 3: Distributions of (a) bulk modulus, (b) shear modulus, (c) Young’s modulus, and (d) Poisson’s ratio, computed as VRH averages from DFT and the four uMLIP models. Each violin plot shows the overall distribution, with blue and red dashed lines marking the median and mean, and short lines denoting the extrema.
As shown in Fig. 3, we compare the distributions of bulk, shear, and Young’s moduli and Poisson’s ratio obtained from DFT and from the four uMLIP models. All models reproduce the qualitative DFT trends, but systematic and model-dependent biases persist in the absolute values. For the bulk modulus, all four models yield mean values close to the DFT reference, indicating that volumetric compressibility is robustly captured across models (DFT: 104.1 GPa; models: 100–115 GPa). In contrast, substantially larger discrepancies are observed for the shear and Young’s moduli: CHGNet systematically underestimates both quantities, whereas MACE and SevenNet consistently overestimate them, with MatterSim providing intermediate values closest to DFT (DFT: 47.3/122.5 GPa; CHGNet: 28.6/77.5 GPa; MACE: 58.2/148.1 GPa; SevenNet: 56.3/143.9 GPa; MatterSim: 50.8/130.5 GPa). An opposite tendency is observed for Poisson’s ratio. CHGNet markedly overestimates the DFT mean, whereas MACE and SevenNet slightly underestimate it; MatterSim again provides the closest agreement with DFT (DFT: 0.291; CHGNet: 0.371; MACE: 0.279; SevenNet: 0.282; MatterSim: 0.294).
@@
To enable quantitative evaluation, we present point-wise comparisons of the primary elastic properties in Fig. 4. For the bulk modulus, SevenNet and MACE exhibit the highest consistency with DFT, achieving correlation coefficients of approximately $R\approx 0.94$ and mean absolute errors (MAE) around 15 GPa, outperforming both CHGNet ($R=0.909$) and MatterSim ($R=0.924$). For the shear modulus, MACE attains the highest correlation ($R=0.896$), followed by SevenNet ($R=0.895$), while MatterSim yields intermediate accuracy ($R=0.847$) and CHGNet remains significantly weaker ($R=0.546$). Regarding Young’s modulus, MatterSim yields the mean closest to DFT, but in this correlation-centric assessment MACE attains the higher correlation ($R=0.901$) than MatterSim ($R=0.860$); SevenNet is lower ($R=0.791$), and CHGNet remains the weakest ($R=0.546$). For Poisson’s ratio, a different trend emerges: MACE and MatterSim achieve significantly higher correlations ($R\approx 0.65$) than CHGNet ($R=0.301$) and SevenNet ($R=0.374$), indicating their robustness in capturing ratio-type properties.
@@
Beyond elasticity, stability classification provides another essential benchmark for evaluating model performance. Fig. 5 compares the stability predictions of the four uMLIPs against DFT references. SevenNet and MACE achieve the highest performance, with accuracies of 98.3% and 98.1%, respectively, and F1 scores approaching 0.99, reflecting well-balanced capability in identifying both stable and unstable materials. MatterSim ranks closely behind, while CHGNet reaches only 93.4% accuracy, significantly lower than the others and showing a higher rate of missed unstable samples.
@@
3.2 Systematic Error Analysis
@@
Fig. 6 presents the relative error distributions of the four uMLIPs with respect to DFT values across various elastic descriptors and the values in parentheses in this paragraph correspond to median relative errors. CHGNet exhibits pronounced systematic deviations across most properties. In bulk modulus predictions, it shows a median error of $2.61\%$, indicating a tendency toward underestimation, whereas MACE and SevenNet display slight overestimations (4.16% and 2.88%, respectively), and MatterSim remains close to zero bias ($1.98\%$). For the shear modulus and Young’s modulus, CHGNet strongly underestimates both ($48.02\%$ and $44.20\%$), in sharp contrast to the overestimations observed for MACE (13.83% and 12.43%) and SevenNet (9.79% and 8.89%), while MatterSim again yields nearly symmetric distributions ($2.12\%$ and $2.24\%$). For Poisson’s ratio, CHGNet systematically overestimates (27.25%), opposite to the mild underestimations of MACE ($4.35\%$) and SevenNet ($3.40\%$), whereas MatterSim remains almost unbiased (0.70%). The bulk/shear ratio further highlights CHGNet’s strong positive bias (77.05%), while the other models show values close to zero. CHGNet also shows especially high variability in anisotropy metrics, suggesting instability in capturing complex anisotropic behavior. For Cauchy pressure, CHGNet exhibits a systematic positive bias, whereas the others lean toward negative deviations. The large deviations in the predicted anisotropy and Cauchy pressure mainly reflect the high sensitivity of these quantities to small differences among elastic constants. In this work, most materials exhibit a relatively small degree of elastic anisotropy, with a DFT average of 1.97. The Cauchy pressure, being a difference quantity defined as $(C_{12}-C_{44})$, has a DFT average value of 17.9 GPa, which is much smaller than the bulk and Young’s modulus that are generally on the order of 100 GPa. Consequently, even moderate relative errors in the stiffness components can lead to large percentage deviations in these derived quantities. Finally, in Debye temperature predictions, CHGNet again underestimates ($25.89\%$), while MACE (6.55%) and SevenNet (4.89%) perform closer to DFT, and MatterSim achieves the most balanced performance ($0.69\%$).
@@
To provide a clearer comparison of overall performance, Fig. 7 summarizes the MAPE values of different models across all elastic properties. It is evident that CHGNet systematically yields the highest error levels, with an average MAPE of 71.8%, underscoring its structural deficiencies in elastic property prediction. In contrast, SevenNet consistently achieves the lowest error, with an average MAPE of only 27.53%, highlighting its superior overall accuracy.
@@
3.3 Fundamental Limitations and the Impact of Fine-Tuning
@@
The limited performance of current uMLIPs in elastic property prediction can be partially attributed to training datasets dominated by equilibrium configurations, which do not provide sufficient coverage of strained states that are important for learning accurate mechanical responses. Fine-tuning [46] provides a practical route to mitigating such systematic biases. To evaluate this strategy, we constructed a targeted fine-tuning dataset comprising $185$ materials with the largest baseline errors, together with DFT-computed energies of their deformed structures, thereby explicitly introducing non-equilibrium information into the training domain. The DFT calculation settings are detailed in the Supporting Information. Fine-tuning on the targeted dataset yields well-controlled training errors for all models. The resulting mean absolute errors (MAEs) for energies, forces, and stresses are summarized in Table 1.
Table 1: Training mean absolute errors (MAEs) obtained after fine-tuning.
Model Energy Force Stress
(meV/atom) (meV/Å) (meV/Å3)
CHGNet 24.089 26.831 1.960
MACE 6.170 26.380 5.670
MatterSim 17.890 9.574 0.836
SevenNet 1.692 5.101 0.528
All four uMLIPs — CHGNet, MACE, MatterSim, and SevenNet — were fine-tuned on this dataset. The updated potentials were then reassessed on the same $185$ materials. Fig. S2 and Fig. S3 display heatmaps of the MAPE before and after fine-tuning, while Fig. S4 and Fig. S5 present the corresponding distributions of relative errors. To quantify these effects, Table 2 summarizes the relative change in MAPE, defined as the difference between the fine-tuned MAPE and the original MAPE, normalized by the original value. CHGNet shows the largest improvement, with an average MAPE reduction of 23.2%, followed by MatterSim at 20.7% and SevenNet at 18.0%. In contrast, MACE’s average MAPE increases by 13.8%, indicating a less favorable response to the added deformation information.
@@
4 Discussion
4.1 Implications for Materials Design Applications
@@
4.2 Future Directions
As demonstrated in our fine-tuning study, incorporating strained configurations into the training process can effectively improve the predictive accuracy of uMLIPs for elastic properties, underscoring the importance of dataset diversity. Future developments should focus on systematically incorporating deformed structures into training datasets, for instance through active learning strategies that aim to improve mechanical property accuracy [47]. For elastic property predictions within specific chemical spaces, constructing domain-specific fine-tuning [28] datasets and adapting pretrained models accordingly could effectively mitigate systematic biases in those regions.
@@ 4.2 Future Directions
5 Conclusions
Our benchmark study establishes the first systematic evaluation framework for applying uMLIPs to elastic property prediction, validated across nearly 11 00011\,000 crystalline materials. The results demonstrate clear differences in model suitability: SevenNet delivers the highest overall accuracy, MatterSim and MACE achieve a favorable balance between accuracy and computational efficiency, while CHGNet, constrained by its architectural design, performs relatively less effectively. These results provide quantitative, evidence-based guidance for selecting suitable uMLIPs in mechanical property calculations, ensuring that model performance is properly matched to application requirements.
@@ Acknowledgements
We acknowledge the Materials Project team for providing comprehensive elastic property data and maintaining the open-access materials database. Parts of the workflow for structure relaxation and elastic property calculations utilized or were adapted from the open-source MatCalc code developed by the Materials Project and Materials Virtual Lab.
@@
All 10 99410\,994 crystal structures and the corresponding DFT-computed elastic property data used in this study are publicly available through the Materials Project database (https://materialsproject.org).
@@
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During manuscript preparation, we utilized GPT-5 LLM to enhance sentence structure, readability, and coherence. We also used Claude 4 LLM for optimizing the Python plotting and simulation scripts. After using these tools, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.